# Model Poisoning Attacks to Federated Learning via Multi-Round Consistency

### v2

@@ 6.3 Impact of FL Settings
(a) M-Krum
@@
(d) N-Bound
@@
Impact of the fraction of fake clients:  Fig. 3 shows the testing error rate of the global model as the fraction of fake clients increases from 1% to 20%. First, we observe that PoisonedFL is more effective, i.e., the testing error rate increases, as more fake clients are injected. This is because the aggregated model update in each training round is more likely to have the sign vector $\bm{s}$ when more fake clients send the malicious model updates with the sign vector $\bm{s}$ to the server. Second, using only 5% of fake clients, our PoisonedFL can already break all FL defenses except Multi-Krum and FLAME. However, existing attacks cannot reach such effectiveness even if 20% of fake clients are injected, except MPAF under Norm Bound defense.
Impact of the participation rate:  In each training round, the server typically selects a fraction (called participation rate) of the clients to participate in training. We use participation rate of 0.1 by default in our experiments. Fig. 6 shows the testing error rates under different defenses and attacks when the participation rate ranges from 0.01 to 1, where FashionMNIST dataset is used. Note that for these experiments, we use FashionMNIST instead of CIFAR-10 dataset, due to limited memory of our GPU server. Moreover, we do not have results for Multi-Krum in Fig. 6 when the participation rate is 1 due to the same reason on insufficient memory. For FLCert, the smallest participation rate is set to 0.03 instead of 0.01, because FLCert divides the clients into 10 groups, and if a participation rate of 0.01 is used, only one client is selected to participate in each training round. Consequently, attacks requiring genuine local models are not applicable. We observe that participation rate almost has no impact on the effectiveness of both existing attacks and our attack. Moreover, our attack consistently outperforms existing attacks across all considered participation rates.
6.4 Impact of Different PoisonedFL Variants
PoisonedFL crafts a malicious model update in each training round as a product of a given sign vector and a dynamically set magnitude vector, which is further decomposed into a product of an unit magnitude vector and a scaling factor. The unit magnitude vector and scaling factor play important roles in a malicious model update. Thus, we study different variants for designing them. We show results on CIFAR-10 dataset for simplicity. Due to limited space, we defer the discussion on impact of the sign vector in Appendix A.2.
@@ 6.4 Impact of Different PoisonedFL Variants
  • Adaptive scaling factor: This variant, adopted by PoisonedFL, leverages how the global model changes in the previous training round to set the scaling factor (i.e., Equation 7) to avoid the malicious model updates are filtered out by the server.
@@
7 Countermeasures and Adaptive Attacks
@@
We extract features for clients based on how PoisonedFL crafts malicious model updates. Depending on which information of the malicious model updates is used to extract features, we consider two variants of GMM, i.e., GMM-Sign and GMM-Magnitude, which extract features based on the magnitudes and signs of the model updates, respectively.
7.1 GMM-Sign
@@
$\begin{split}\bm{s}^{t}_{i}[j]&=\begin{cases}-\bm{s}{[j]}&\text{with probability }\alpha\\ \bm{s}{[j]}&\text{with probability }1-\alpha\end{cases},\\ \bm{k}^{t}_{i}[j]&=\begin{cases}\epsilon&\text{if }\bm{s}^{t}_{i}[j]=-\bm{s}{[j]}\\ \bm{k}^{t}{[j]}&\text{otherwise }\end{cases},\end{split}$ (9)
@@
(a) Median
(b) TrMean
(c) Norm Bound
@@
Experimental results:  Fig. 8 shows the detection accuracy of GMM-Sign and the testing error rates as the flipping rate $\alpha$ varies when the server uses different aggregation rules, where the dataset is CIFAR-10 and $\epsilon=1e{-6}$. We observe that GMM-Sign can accurately detect the fake clients when $\alpha$ is small. However, the detection accuracy reduces to 0 and PoisonedFL-Adapt still compromises the learnt global models when $\alpha$ is larger than a threshold, e.g., 0.08.
7.2 GMM-Magnitude
Feature:  PoisonedFL uses the same malicious model update (i.e., $\bm{g}_{i}^{t}=\bm{k}^{t}\odot\bm{s}$) on all fake clients in training round $t$. Therefore, the server can extract features for the clients based on the similarity between their model updates. Specifically, for each client $i$ in training round $t$, we propose to calculate a feature $x_{i}^{t}$ as the sum of the average square distance between the client’s model update and its closest $m-1$ model updates in the recent $N$ training rounds. Formally, we have:
@@ 7.2 GMM-Magnitude
where $\mathbb{N}_{i}^{t^{\prime}}$ is the set of $m-1$ clients whose model updates are the closest to that of client $i$ in round $t^{\prime}$ and $m$ is the number of fake clients. Note that we give advantages to the defense by assuming that the number of fake clients is known to the server. In a round, the server divides the clients into two clusters via GMM based on the features. The cluster with a smaller mean value is detected as fake clients. The mean value is zero for PoisonedFL since the malicious model updates are the same for the fake clients in a round.
@@
$\bm{k}^{t}_{i}=\bm{k}^{t}+\gamma\cdot\frac{||\bm{k}^{t}||}{||\bm{\epsilon}_{i}||}\cdot\bm{\epsilon}_{i},$ (11)
where $\bm{k}^{t}$ is the magnitude vector crafted by PoisonedFL, $\bm{k}^{t}_{i}$ is the magnitude vector for fake client $i$ crafted by PoisonedFL-Adapt, $\bm{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the random noise sampled from the standard Gaussian distribution, $\frac{||\bm{k}^{t}||}{||\bm{\epsilon}_{i}||}$ normalizes the random noise to have the same magnitude as $\bm{k}^{t}$, and $\gamma$ is a scaling factor for the random noise. The noise makes the malicious model updates of the fake clients less similar to each other and thus evade detection.
@@
7.3 Normalization-based Defense
PoisonedFL aims to increase the magnitude of the total aggregated model update, i.e., $\left\|\bm{w}^{T}-\bm{w}^{0}\right\|$. Therefore, one tailored defense is to normalize the total aggregated model update. Specifically, the server can normalize $\bm{w}^{T}-\bm{w}^{0}$ to have a predefined $\ell_{2}$ norm $b$ and add the normalized total aggregated model update to the initial global model $\bm{w}^{0}$ to get a new global model, i.e., $\bm{w}^{T}\leftarrow\bm{w}^{0}+b\frac{\bm{w}^{T}-\bm{w}^{0}}{\left\|\bm{w}^{T}-\bm{w}^{0}\right\|}$. Fig. 11 in Appendix shows the testing error rate of the final global model under PoisonedFL as the predefined norm $b$ varies. Our results show that the testing error rate can be decreased by this defense but is still high. This is because the update direction of the total aggregated model update is randomly picked by the attacker, which normalization cannot change.
8 Discussion and Limitations
Using synthetic data:  In our previous experiments, when applying attacks that require genuine clients’ local models to craft malicious model updates, we assume the attacker has access to the local models on all genuine clients. This assumption is strong and not realistic. Therefore, we also explore that an attacker uses the global models to reconstruct synthetic data, train local models using the synthetic data, and then craft malicious model updates with them.
Reconstructing training data from a model is known as model inversion [39]. Previous work [40] shows that the global-model trajectory in multiple training rounds (especially early ones) of FL can be used to reconstruct synthetic data that is similar to the clients’ genuine local training data. We follow the setting and generation strategy in [40], which is the state-of-the-art reconstruction method in FL, to generate synthetic data for MNIST and FashionMNIST. We use MNIST and FashionMNIST because they are simpler datasets and easier to reconstruct, giving advantages to these attacks. Specifically, we assume no attacks in the first $40$ rounds and use the corresponding clean global models to reconstruct synthetic data. Table 2 shows the testing error rates of the classifiers trained on the synthetic data and evaluated on the genuine testing datasets. The results show that the classifiers have decent testing error rates and thus the synthetic data mimic the genuine data well, which is consistent with prior work [40].

### v3

@@ 5.1.3 Defenses and Compared Attacks
5.1.4 FL Settings
By default, we assume $1,200$ genuine clients and 20% of fake clients, i.e., the number of fake clients is 20% of the number of genuine clients. In each training round, the server selects a random subset of 10% of clients to participate in training. A genuine client trains its local model for 1 local epoch in a training round using SGD. Due to limited space, more FL settings for the five datasets are detailed in Appendix B.4. We also study the impact of different FL settings (fraction of fake clients, non-IID degree, local training epochs, and participation rate) in Section 5.3. Our PoisonedFL has parameters $c^{t}$, $e$, and $\beta$. We set $c^{0}=8$, $e=50$, and $\beta=0.7$, and we do not decrease $c^{t}$ to be smaller than $0.5$ since a too small $c^{t}$ is ineffective. We conduct our experiments using 16 GeForce RTX 3090 GPUs.
5.2 Main Results
Table 1 shows the testing error rates of the learnt global models under different attacks and defenses for the five datasets. “No Attack” means no fake clients are injected.
PoisonedFL breaks state-of-the-art FL defenses:  PoisonedFL substantially increases the testing error rates of the global models learnt by FedAvg and state-of-the-art defenses, compared to No Attack. Specifically, PoisonedFL increases the testing error rate by 15.84 to 90.67 across the five datasets and nine FL methods. Moreover, in 29 out of the 45 cases, PoisonedFL makes the global model nearly random guessing or worse. Our results show that existing FL defenses are not robust even if an attacker has no knowledge about the defense nor the genuine clients’ local training data or models. We detail why PoisonedFL breaks FL defenses in Appendix C.1.
PoisonedFL outperforms existing attacks:  For the non-robust FedAvg, both PoisonedFL and some existing attacks can reduce the global models to random guessing or worse. However, for the five datasets and eight FL defenses, our PoisonedFL achieves a much larger testing error rate than existing attacks in 32 out of the 40 cases with the gap ranging from 2.11 to 84.33. In six of the remaining cases, both our PoisonedFL and some existing attacks make the learnt global models nearly random guessing or worse. Our PoisonedFL achieves a smaller testing error rate than some existing attacks only in two cases, where PoisonedFL still achieves large testing error rates even in these rare cases.
Among the three attacks Random, MPAF, and PoisonedFL, which do not require knowledge about the local training data/models on the genuine clients nor the defense, Random is ineffective in most cases when a defense is deployed; while MPAF is effective to some extent at attacking some defenses, e.g., TrMean and Norm Bound, but ineffective at attacking other defenses, e.g., Multi-Krum, FLTrust, and FLDetector. PoisonedFL substantially outperforms Random and MPAF when a defense is deployed in all cases except (CIFAR10, Norm Bound), for which both MPAF and PoisonedFL make the global models random guessing. For instance, on MNIST dataset, PoisonedFL achieves 59.24 to 85.87 higher testing error rates than MPAF at attacking a defense. We detail why PoisonedFL outperforms existing attacks in Appendix C.1.
5.3 Impact of FL Settings
We study the impact of different FL settings, including the fraction of fake clients, degree of non-IID, the number of local training epochs, and participation rate, on the effectiveness of existing attacks and our attack. Unless otherwise mentioned, we show experimental results on CIFAR-10 dataset for simplicity, but we have similar observations on other datasets. Moreover, when studying the impact of one parameter, we use the default settings for other parameters discussed in Section 5.1.4.
Impact of the fraction of fake clients:  Fig. 3 shows the testing error rate of the global model as the fraction of fake clients increases from 1% to 20%. First, we observe that our PoisonedFL is more effective, i.e., the testing error rate increases, as more fake clients are injected. This is because the aggregated model update in each training round is more likely to have the sign vector $\bm{s}$ when more fake clients send the malicious model updates with the sign vector $\bm{s}$ to the server. Second, using only 5% of fake clients, our PoisonedFL can already break all FL defenses except Multi-Krum and FLAME. However, existing attacks cannot reach such effectiveness even if 20% of fake clients are injected, except MPAF under Norm Bound defense.
Due to limited space, we show the impact of participation rate and ablation study on variants of PoisonedFL in Appendix C.2, C.3, and C.4.
6 Discussion and Limitations
Countermeasures and adaptive attacks:  If the attack is known, we propose tailored defenses. Specifically, we study a new defense that normalizes the total aggregated model update since PoisonedFL aims to increase its magnitude. We also explore GMM-Sign and GMM-Magnitude to detect fake clients based on how PoisonedFL crafts the malicious model updates. We show that the normalization-based defense has limited effectiveness at mitigating PoisonedFL and we can still adapt PoisonedFL to break the detection based defenses. Due to limited space, the detailed methods and evaluation are deferred to Appendix D.1.
Further discussion on the potential application of synthetic data based attacks and limitations in cross-silo FL is specified in Appendix D.2 and D.3, respectively.
7 Conclusion
Supplementary Material
Appendix A Motivation
We first introduce our key observation about why existing model poisoning attacks achieve suboptimal attack effectiveness, which motivates the design of our PoisonedFL in the next section. We show experimental results on the MNIST dataset with the default parameter settings described in Section 5.1. We assume the server uses Trimmed Mean [39] as the aggregation rule. Moreover, we assume the ratio of fake clients to genuine clients is $20\%$. We consider MPAF [13], an attack not requiring genuine clients’ information, and Fang [17], a representative attack that requires genuine clients’ information. To give advantages to Fang, we assume the attacker knows the local models of all participating genuine clients in each training round when crafting malicious model updates on fake clients.
In each training round, these attacks craft a malicious model update for each fake client selected by the server to participate in training. A malicious model update $\bm{g}_{i}^{t}$ of fake client $i$ in training round $t$ can be decomposed into a dimension-wise product of a sign vector $\bm{s}_{i}^{t}$ and a magnitude vector $|\bm{g}_{i}^{t}|$, where $|\cdot|$ means dimension-wise absolute value of a vector. Each dimension of $\bm{s}_{i}^{t}$ is either +1 or -1, while all dimensions of $|\bm{g}_{i}^{t}|$ are non-negative. A dimension with a +1 (or -1) sign aims to increase (or decrease) the corresponding dimension/parameter of the global model. For a fake client, we say a dimension is flipped in training round $t$ if its sign in the malicious model update is flipped, compared to that in the previous training round. A flipped dimension in a training round $t$ may cancel the attack effect of the previous training round t1t-1. This is because the malicious model updates aim to increase the dimension of the global model in one training round but decrease it in the other. For each training round $t$, we define flipping rate as the fraction of flipped dimensions of a malicious model update averaged over the fake clients.
Fig. 1(a) shows the testing error rate of the global model as a function of training round when no attack, MPAF, or Fang is used; while Fig. 1(b) shows the flipping rate as a function of training round when MPAF or Fang is used to craft malicious model updates on the fake clients. We observe that the flipping rate of MPAF is always larger than 35% and that of Fang is larger than 40% in the last around 1,500 training rounds, leading to self-cancellation of attack effect in many training rounds. For instance, in Fang attack, the flipping rate is around 10% in nearly the first 1,000 training rounds, i.e., the malicious model updates maintain some degree of “consistency”. As a result, the global model under Fang attack has a large testing error rate in those training rounds as shown in Fig. 1(a). However, the flipping rate rapidly increases as the training proceeds, eventually leading to self-cancellation of attack effect and an accurate global model.
Appendix B Additional Details of Experimental Setup
Table 2: Model architecture for MNIST and FashionMNIST.
Layer Type Size
Convolution + ReLU 3x3x30
Max Pooling 2x2
Convolution + ReLU 3x3x50
Max Pooling 2x2
Fully Connected + ReLU 100
Softmax 10
Table 3: Model architecture for CIFAR-10.
Layer Type Size
Convolution + ReLU 3x3x32
Max Pooling 2x2
Convolution + ReLU 3x3x64
Max Pooling 2x2
Fully Connected + ReLU 512
Softmax 10
Table 4: Model architecture for FEMNIST.
Layer Type Size
Convolution + ReLU 3x3x30
Max Pooling 2x2
Convolution + ReLU 3x3x50
Max Pooling 2x2
Fully Connected + ReLU 200
Softmax 62
B.1 Datasets
We use the following five datasets from different domains.
B.2 Defenses
Multi-Krum [11]:  Multi-Krum uses an iterative method to select a subset of clients’ model updates. In each step, it selects the model update that has the smallest sum of Euclidean distance to its n2n-2 neighbors, where $n$ is the number of genuine clients. This selection process continues until $n$ clients are chosen. Finally, the server computes the average of the $n$ selected model updates as the aggregated model update.
Trimmed Mean (TrMean) [39]:  For each dimension, the server first removes the largest $m$ values and smallest $m$ values, and then computes the average of the remaining nmn-m values, where $m$ is the number of fake clients.
B.3 Compared Attacks
We compare our PoisonedFL with seven attacks, including five attacks (Fang [17], Opt. Fang [32], LIE [10], Min-Max [32], and Min-Sum [32]) that require genuine clients’ information and two attacks (Random [13] and MPAF [13]) that do not require such information. Note that when applying the first category of attacks to craft the malicious model updates on fake clients, we assume that the attacker has access to the model updates of all genuine clients and the aggregation rule used by the server. In Section D, we also study the scenarios where an attacker uses the global models to reconstruct synthetic data on the fake clients and uses the local models trained on them to perform attacks. We do not make these assumptions for Random, MPAF, and our PoisonedFL. Therefore, we give advantages to the first category of attacks.
MPAF [13]:  In MPAF attack, during training round $t$, fake client $i\in[n+1,n+m]$ crafts its malicious model update as $\bm{g}_{i}^{t}=\lambda\cdot(\bm{w}^{\prime}-\bm{w}^{t-1}$), where $\bm{w}^{\prime}$ is an attacker-chosen random model, $\bm{w}^{t-1}$ is the global model in the previous training round t1t-1, and $\lambda$ is a scaling factor. Following prior work [13], we set $\lambda=1e6$ in our experiments.
B.4 FL Settings
We primarily establish the FL setting in a cross-device scenario, which is more feasible for model poisoning attacks, as discussed in Section D. Therefore, we assume $1,200$ genuine clients and $20\%$ fake clients. The learning rates for the five datasets are selected within the range of $0.01$ to $0.1$ for optimal training effectiveness. Specifically, the learning rates for MNIST, FashionMNIST, and FEMNIST are set to $0.01$, while CIFAR-10 is set to $0.03$, and Purchase is set to $0.1$. The number of training rounds is configured to ensure full convergence under various scenarios: $6,000$ for MNIST, FashionMNIST, and FEMNIST, and $10,000$ for CIFAR-10 and FEMNIST. The batch size is set to $32$ for MNIST, FashionMNIST, CIFAR-10, and FEMNIST, and to $128$ for Purchase, based on the number of local data samples.
Appendix C Additional Experimental Results
C.1 Analyzing the Effectiveness of PoisonedFL
Why PoisonedFL breaks FL defenses:  State-of-the-art defenses leverage filtering and clipping to reduce the impact of malicious model updates. They can weaken the attack effect in individual training rounds. However, since PoisonedFL crafts malicious model updates that are consistent across training rounds and avoid being entirely filtered out by dynamically adjusting attack magnitudes, its cumulative attack effect still substantially moves the global model towards the designated random update direction $\bm{s}$, leading to a large testing error rate. For instance, Fig. 7 in Appendix shows that a large fraction of the dimensions of the total aggregated model updates have signs matching with $\bm{s}$ and Fig. 8 in Appendix shows that the magnitude of the total aggregated model updates becomes very large, after our attack effect accumulates over multiple training rounds, for all defenses.
Specifically, for Median and TrMean, even though the malicious model updates are not necessarily selected, their consistent nature over multiple training rounds biases the aggregation of each dimension towards the designated random update direction. For Norm Bound, even if our attack effect is weakened in each training round, its cumulative effect over multiple training rounds remains pronounced with a significant magnitude for the total aggregated model update. For filtering-involved defenses based on Euclidean distance or cosine similarity, such as Multi-Krum, FLTrust, and FLAME, PoisonedFL maintains effectiveness because the malicious model updates are not entirely filtered out in many training rounds due to dynamic magnitude adjustment. For instance, Table 5 in Appendix shows that a large fraction of or all malicious model updates are not filtered out by FLAME and Multi-Krum, while Table 6 shows similar results for FLTrust. FLCert is vulnerable to PoisonedFL since the global models for ensembling are learnt by existing aggregation rules, which are vulnerable to PoisonedFL. The detection mechanism FLDetector relies on inconsistent model updates of malicious clients. Since the malicious model updates are consistent in PoisonedFL, FLDetector falsely classifies all malicious clients as benign across all datasets.
Why PoisonedFL outperforms existing attacks:  PoisonedFL outperforms existing attacks because their crafted malicious model updates are inconsistent across training rounds, leading to self-cancellation of attack effect. In contrast, PoisonedFL enforces the malicious model updates to be consistent (i.e., have the same sign vector) across training rounds, making the aggregated model updates add up in the global model across training rounds. As a result, the final learnt global model has a very large magnitude, leading to a large testing error rate. For instance, Fig. 8 in Appendix shows the magnitude (measured by $\ell_{2}$ norm) of the total aggregated model updates as a function of training round for different attacks on the Purchase dataset. Our results show that the total aggregated model updates under existing attacks have small magnitudes while those under our PoisonedFL attack have much larger magnitudes, which is caused by the inconsistent malicious model updates in existing attacks vs. consistent malicious model updates in our attack.
Figure 7: Fraction of dimensions of the aggregated model update whose signs match with the sign vector $\bm{s}$ as a function of training round under PoisonedFL. The dataset is Purchase.
(a) Multi-Krum
@@ C.1 Analyzing the Effectiveness of PoisonedFL
(d) Norm Bound
(e) FLTrust
(f) FLAME
Figure 8: $\ell_{2}$ norm of the total aggregated model updates $\left\|\sum_{t^{\prime}=1}^{t}\bm{g}^{t^{\prime}}\right\|$ (i.e., $\left\|\bm{w}^{t}-\bm{w}^{0}\right\|$) as a function of training round $t$ for different attacks on the Purchase dataset.
Multi-Krum FLAME
[…]
Malicious Benign
[…]
(a) Multi-Krum
(b) Median
(c) TrMean
(d) Norm Bound
@@
C.2 Impact of the Participation Rate
In each training round, the server typically selects a fraction (called participation rate) of the clients to participate in training. We use participation rate of 0.1 by default in our experiments. Fig. 9 shows the testing error rates under different defenses and attacks when the participation rate ranges from 0.01 to 1, where FashionMNIST dataset is used. Note that for these experiments, we use FashionMNIST instead of CIFAR-10 dataset, due to limited memory of our GPU server. Moreover, we do not have results for Multi-Krum in Fig. 9 when the participation rate is 1 due to the same reason on insufficient memory. For FLCert, the smallest participation rate is set to 0.03 instead of 0.01, because FLCert divides the clients into 10 groups, and if a participation rate of 0.01 is used, only one client is selected to participate in each training round. Consequently, attacks requiring genuine local models are not applicable. We observe that participation rate almost has no impact on the effectiveness of both existing attacks and our attack. Moreover, our attack consistently outperforms existing attacks across all considered participation rates.
C.3 Impact of Different PoisonedFL Variants
PoisonedFL crafts a malicious model update in each training round as a product of a given sign vector and a dynamically set magnitude vector, which is further decomposed into a product of an unit magnitude vector and a scaling factor. The unit magnitude vector and scaling factor play important roles in a malicious model update. Thus, we study different variants for designing them. We show results on CIFAR-10 dataset for simplicity.
@@ C.3 Impact of Different PoisonedFL Variants
  • Adaptive scaling factor: This variant, adopted by PoisonedFL, leverages how the global model changes in the previous training round to set the scaling factor (i.e., Equation 7) to avoid that the malicious model updates are filtered out by the server.
@@
C.4 Impact of the Sign Vector
seed=1 seed=2 seed=3 seed=4 seed=5 seed=6
Norm Bound 90.00 90.00 89.95 90.00 88.94 90.00
FLTrust 90.00 88.15 89.09 90.00 90.00 90.00
FLAME 80.79 86.59 84.43 81.39 80.68 82.37
We further analyze the impact of the random sign vectors. We experiment with PoisonedFL on CIFAR-10 with $6$ random seeds (11-66) for various random sign vectors and defenses including TrMean, FLTrust, and FLAME, as shown in Table 7. The result aligns with our intuition that these random sign vectors do not affect the attack performance and PoisonedFL can successfully break the defenses with these $6$ different random sign vectors. The reason is that in a continuous attack with the same random direction, the model inevitably updates substantially in a random direction in an uncontrolled way, and therefore leads to significantly degraded accuracy.
Appendix D Discussion and Limitations
D.1 Countermeasures and Adaptive Attacks
(a) Median
(b) TrMean
(c) Norm Bound
@@ D.1 Countermeasures and Adaptive Attacks
We extract features for clients based on how PoisonedFL crafts malicious model updates. Depending on which information of the malicious model updates is used to extract features, we consider two variants of GMM, i.e., GMM-Magnitude and GMM-Sign, which extract features based on the magnitudes and signs of the model updates, respectively. Due to limited space, we discuss GMM-Magnitude in Appendix D.1.2.
D.1.1 GMM-Sign
@@
[…]
@@
Experimental results:  Fig. 11 shows the detection accuracy of GMM-Sign and the testing error rates as the flipping rate $\alpha$ varies when the server uses different aggregation rules, where the dataset is CIFAR-10 and $\epsilon=1e{-6}$. We observe that GMM-Sign can accurately detect the fake clients when $\alpha$ is small. However, the detection accuracy reduces to 0 and PoisonedFL-Adapt still makes the learnt global models random guessing when $\alpha$ is larger than a threshold, e.g., 0.08.
@@
D.1.2 GMM-Magnitude
Feature:  PoisonedFL uses the same malicious model update (i.e., $\bm{g}_{i}^{t}=\bm{k}^{t}\odot\bm{s}$) on all fake clients in training round $t$. Therefore, the server can extract features for the clients based on the similarity between their model updates. Specifically, for each client $i$ in training round $t$, we propose to calculate a feature $x_{i}^{t}$ as the sum of the average square distance between the client’s model update and its closest m1m-1 model updates in the recent $N$ training rounds. Formally, we have:
@@ D.1.2 GMM-Magnitude
where $\mathbb{N}_{i}^{t^{\prime}}$ is the set of m1m-1 clients whose model updates are the closest to that of client $i$ in training round $t^{\prime}$ and $m$ is the number of fake clients. Note that we give advantages to the defense by assuming that the number of fake clients is known to the server. In a training round, the server divides the clients into two clusters via GMM based on the features. The cluster with a smaller mean value is detected as fake clients. The mean value is exactly zero for PoisonedFL since the malicious model updates are the same for the fake clients in each training round.
@@
[…]
@@
D.1.3 Normalization-based Defense
PoisonedFL aims to increase the magnitude of the total aggregated model update, i.e., $\left\|\bm{w}^{T}-\bm{w}^{0}\right\|$. Therefore, one tailored defense is to normalize the total aggregated model update. Specifically, the server can normalize $\bm{w}^{T}-\bm{w}^{0}$ to have a predefined $\ell_{2}$ norm $b$ and add the normalized total aggregated model update to the initial global model $\bm{w}^{0}$ to get a new global model, i.e., $\bm{w}^{T}\leftarrow\bm{w}^{0}+b\frac{\bm{w}^{T}-\bm{w}^{0}}{\left\|\bm{w}^{T}-\bm{w}^{0}\right\|}$. Fig. 13 shows the testing error rate of the final global model under PoisonedFL as the predefined norm $b$ varies. Our results show that the testing error rate can be decreased by this defense but is still high. This is because the update direction of the total aggregated model update is randomly picked by the attacker, which normalization cannot change.
D.2 Using Synthetic Data
In our previous experiments, when applying attacks that require genuine clients’ local models to craft malicious model updates on fake clients, we assume the attacker has access to the local models on all genuine clients. This assumption is strong and not realistic. Therefore, we also explore that an attacker uses the global models to reconstruct synthetic data, train local models using the synthetic data, and then craft malicious model updates based on them.
Reconstructing training data from a model is known as model inversion [20]. Previous work [30] shows that the global-model trajectory in multiple training rounds (especially early ones) of FL can be used to reconstruct synthetic data that is similar to the clients’ genuine local training data. We follow the setting and generation strategy in [30], which is the state-of-the-art reconstruction method in FL, to generate synthetic data for MNIST and FashionMNIST. We use MNIST and FashionMNIST because they are simpler datasets and easier to reconstruct, giving advantages to these attacks. Specifically, we assume no attacks in the first 40 training rounds and use the corresponding clean global models to reconstruct synthetic data. Table 8 shows the testing error rates of the classifiers trained on the synthetic data and evaluated on the genuine testing datasets of MNIST and FashionMNIST. The results show that the classifiers have decent testing error rates and thus the synthetic data mimic the genuine data well. These results are also consistent with prior work [30].